OpenStudy (anonymous):

(3-4 square root7)(3-4 square root7)

OpenStudy (anonymous):



OpenStudy (anonymous):

I need the steps please.

OpenStudy (anonymous):

\[(3-4\sqrt{7})^2\] use \[(a-b)^2 = a^2 - 2ab + b^2\]